OpenStudy (anonymous):

g(x)=5x^2+3x-2 determine the domain and range

OpenStudy (tkhunny):

What say you? Have you any reason to believe the Domain is restricted? Have you seen "-b/2a" in relation to minimum values or the Range?

OpenStudy (anonymous):

ahm, think the domain is set of real numbers,, i do not know how to do the range

OpenStudy (tkhunny):

Can you "Complete the Square"?

OpenStudy (anonymous):

huh?? i really can't understand you.

OpenStudy (anonymous):

it said that in our notes is that range is y ≥ 4ac-b^2/4a, if a.0 and y≤0 4ac-b^2/4a , if a<0

OpenStudy (tkhunny):

I'll take that as a "no". Our task is to find the minimum or maximum value of g(x). You're saying you have no way of finding this value? 1) That 5 out front is positive. This is a parabola that opens UP, therefore we will not be finding a maximum value. We well need to search for a minimum value. 2) Completing the Square is a great way to find such a value. However, completing the square ALWAYS results in the solution x = -b/(2a) for a general quadratic. Let's try it. From g(x), we have a = 5, b = 3, c -2 -- like you would utilize the quadratic formula, right? So x = -3/[2*5] = -3/10. Please evaluate g(-3/10) and see if you can produce a value of g(x) that is less than g(-3/10). Note: When I am using the right language for things you should know, you not understanding me is not encouraging. Why have you not been given the tools to solve the problems you have ben given? Have you been sleeping in class? Are you reading your course materials? Gotta step up your game a bit, I would think.